OpenStudy (anonymous):

Evaluate C(n,n). ans: 0, 1, or n?

OpenStudy (campbell_st):

this looks like a combination question so you need to know 0! = 1 so using the combination formula \[\frac{n!}{n!(n - n)!} = \frac{n!}{n! \times 1}\] I'll let finish off to find the value.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i'm lost, so can u just tell me?

OpenStudy (anonymous):

You should memorize that formula. It will be on the test.

OpenStudy (anonymous):

hahh will do